Consider the following hypothesis test: Ho: u <= 12, Ha: u > 12. A sample of 25 provided a sample mean of 14 and a sample standard deviation of 4.32. a) test statistic? b) p- value? c) conclusion with level of significance at .05?

Question 2 options:

2.31, .01-.025, reject Ho

2.31, .01-.025, fail to reject Ho

3.2, .02-.05, reject Ho

3.2, .02-.05, fail to reject Ho
Answer #1

Test statistics

t = ( - ) / (S / sqrt(n ) )

= ( 14 - 12) / ( 4.32 / sqrt(25) )

= 2.31

df = n - 1 = 25 - 1 = 24

From T table,

Probability for the test statistics 2.31 and df of 24 is approximately, 0.015

P-value lies between 0.01 - 0.025

Since p-value < 0.05 significance level, Reject H0.

Answer: 2.31 , 0.01 - 0.025 , Reject H0.